# SLR: Automated Synthesis for Scalable Logical Reasoning

### v4

@@ Appendix C Compute Costs and Model Pricing
Compute costs are reported as the total USD cost to run all prompts, based on publicly listed API prices as of 01.05.2025. Pricing ignores server-side token caching, as actual cache hit counts are unavailable. Table 5 summarizes per-model cost rates and API sources.
⟦t⟧ Model Model Tag Input Input(Cached) Output API gpt-4.1 gpt-4.1-2025-04-14 2.00 0.50 8.00 OpenAI gpt-4.1-mini gpt-4.1-mini-2025-04-14 0.40 0.10 1.60 OpenAI gpt-4.1-nano gpt-4.1-nano-2025-04-14 0.10 0.025 0.40 OpenAI gpt-4.5-preview gpt-4.5-preview-2025-02-27 75.00 37.50 150.00 OpenAI gpt-4o gpt-4o-2024-08-06 2.50 1.25 10.00 OpenAI gpt-4o-mini gpt-4o-mini-2024-07-18 0.15 0.075 0.60 OpenAI o1 o1-2024-12-17 15.00 7.50 60.00 OpenAI o1-pro o1-pro-2025-03-19 150.00 600.00 OpenAI o3 o3-2025-04-16 10.00 2.50 40.00 OpenAI o4-mini-low o4-mini-2025-04-16 1.10 0.275 4.40 OpenAI o4-mini o4-mini-2025-04-16 1.10 0.275 4.40 OpenAI o4-mini-high o4-mini-2025-04-16 1.10 0.275 4.40 OpenAI o3-mini o3-mini-2025-01-31 1.10 0.55 4.40 OpenAI o1-mini o1-mini-2024-09-12 1.10 0.55 4.40 OpenAI DeepSeek-R1 DeepSeek-R1-Distill-Llama-70B 0.10 0.40 OpenRouter Internlm2-20b Internlm2-20b 0.15 0.20 Estimated Llama-3.2-3B Llama-3.2-3B-Instruct 0.015 0.025 OpenRouter Llama-3.1-8B Llama-3.1-8B-Instruct 0.02 0.03 OpenRouter Llama-3.3-70B Llama-3.3-70B-Instruct 0.10 0.25 OpenRouter Mixtral-8x7B Mixtral-8x7B-Instruct-v0.1 0.24 0.24 OpenRouter Qwen2-57B-A14B Qwen2-57B-A14B-Instruct 0.70 0.70 Estimated QwQ-32B QwQ-32B-Preview 0.15 0.20 OpenRouter CodeLlama-34b CodeLlama-34b-Instruct-hf 0.776 0.776 TogetherAI MetaTuned Llama Llama-3.1-8B-Tuned-FFT 0.02 0.03 OpenRouter MetaTuned Llama LoRA Llama-3.1-8B-Tuned-LoRA 0.02 0.03 OpenRouter
Table 5: Model Pricing (\$ per 1M tokens). API rates as of 01.05.2025.
Appendix D Example: Task Synthesis Process
@@ 3. Instruction Prompt (example formats):
    % Your task is to find a rule "is_red_train(T) :-" that solves the bk.
@@
    % Your task is to find a rule "is_red_train(T) :-" that solves the bk.
@@
Appendix E LLM Guided Rule Generation
1. Conjunction with Existential Quantification: There exists a red short car
There is at least one car that is both short and red.
1eastbound(Train) :-
2 has_car(Train, Car),
3 car_color(Car, red),
4 car_len(Car, short).
2. Disjunction: Some car is white or yellow.
At least one car is either white or yellow.
1eastbound(Train) :-
2 has_car(Train, Car),
3 (car_color(Car, white) ; car_color(Car, yellow)).
3. Negation: The train does not contain any red cars
No car on the train is red.
1eastbound(Train) :-
2 \+ (has_car(Train, Car), car_color(Car, red)).
4. Inequality/Distinctness: Two cars must have different colors
There are at least two cars on the train with different colors.
1eastbound(Train) :-
2 has_car(Train, CarA),
3 has_car(Train, CarB),
4 CarA \= CarB,
5 car_color(CarA, Color1),
6 car_color(CarB, Color2),
7 Color1 \= Color2.
5. Aggregation/Counting: There are more green cars than yellow cars
The train contains more green cars than yellow cars.
1eastbound(Train) :-
4 length(Greens, G),
5 length(Yellows, Y),
6 G > Y.
6. Mutual Exclusion: Only one car is yellow; all others are not yellow
There is exactly one yellow car; all others are not yellow.
1eastbound(Train) :-
3 forall(
4 (has_car(Train, Car), Car \= YellowCar),
5 (car_color(Car, NotYellow),
6 NotYellow \= yellow)
7 ).
7. Uniqueness: No two cars have the same color
All cars have unique colors.
1eastbound(Train) :-
3 sort(Colors, UniqueColors),
4 length(Colors, N),
5 length(UniqueColors, N).
8. No-Other/Uniqueness: Only two cars in the train
Only two cars are present in the train.
1eastbound(Train) :-
2 findall(Car, has_car(Train, Car), Cars),
3 length(Cars, 2).
9. Universal Quantification: Every full-wall car is long
All cars with a full wall must be long.
1eastbound(Train) :-
2 forall(
3 (has_car(Train, Car), has_wall(Car, full)),
4 car_len(Car, long)
5 ).
10. Conditional Implication: All long cars are either red or blue
Every long car is either red or blue.
1eastbound(Train) :-
2 forall(
3 (has_car(Train, Car), car_len(Car, long)),
4 (car_color(Car, Color), (Color = red ; Color = blue))
5 ).
11. Conditional Aggregation: All long cars are either red or blue
Every long car is either red or blue.
1eastbound(Train) :-
2 forall(
3 (has_car(Train, Car), car_len(Car, long)),
4 (car_color(Car, Color), (Color = red ; Color = blue))
5 ).
12. Pattern Matching: All full-wall cars are white
Every full-wall car is white.
1eastbound(Train) :-
2 forall(
3 (has_car(Train, Car), has_wall(Car, full)),
4 car_color(Car, white)
5 ).
13. Symmetry: Two cars are neighbors with same color
CarA and CarB are neighbors on the train and have the same color.
1eastbound(Train) :-
2 has_car(Train, CarA),
3 has_car(Train, CarB),
4 CarA \= CarB,
5 car_num(CarA, N1),
6 car_num(CarB, N2),
7 (N2 =:= N1 + 1 ; N2 =:= N1 - 1),
8 car_color(CarA, Color),
9 car_color(CarB, Color).
14. Combinatorial Group: Exactly two short yellow cars
There are exactly two yellow cars, and both are short.
1eastbound(Train) :-
3 length(L, 2).
15. Recursion: At least one long car in the train
The train has at least one long car.
1eastbound([Car|Cars]) :-
2 car_len(Car, long)
3 ;
4 eastbound(Cars).
16. Existence of a Structure (Sublist Pattern Matching)
Exists three cars in sequence: Num, Num+1, Num+2, matching pattern.
1eastbound(Train) :-
2 has_car(Train, Car1), car_num(Car1, N),
3 car_len(Car1, short),
4 N2 is N+1, N3 is N+2,
5 has_car(Train, Car2), car_num(Car2, N2), car_len(Car2, long),
6 has_car(Train, Car3), car_num(Car3, N3), car_len(Car3, short).
17. Min/Max and Extremal Values
A short car followed by a long car followed by a short car, anywhere in the train.
1eastbound(Train) :-
2 findall(N, (has_car(Train, Car), car_num(Car, N)), Numbers),
3 max_list(Numbers, Max),
4 has_car(Train, LastCar),
5 car_num(LastCar, Max),
6 car_color(LastCar, white).
18. Subset/Superset Constraints
All full-wall cars are among the first three cars.
1eastbound(Train) :-
2 forall(
3 (has_car(Train, Car), has_wall(Car, full)),
4 (car_num(Car, N), N =< 3)
5 ).
19. Projection/Aggregation Over Multiple Properties
All pairs of cars have different (color, length) tuples.
1eastbound(Train) :-
2 has_car(Train, CarA), has_car(Train, CarB), CarA \= CarB,
3 car_color(CarA, ColA), car_len(CarA, LenA),
4 car_color(CarB, ColB), car_len(CarB, LenB),
5 (ColA \= ColB ; LenA \= LenB).
20. All-Different on Multiple Attributes
Enforce all car colors are different, AND all car numbers are different (car numbers are unique by assumption, but see structure).
1eastbound(Train) :-
3 sort(Colors, UniqueColors),
4 length(Colors, N), length(UniqueColors, N).
Appendix F First-Order Logic Details
Appendix G Training and Evaluation Details
G.1 Training Setup
For curriculum learning experiments on SLR-Bench, we fine-tune the Llama-3.1-8B-Instruct model using supervised fine-tuning (SFT) with LoRA adapters using LLaMA-Factory zheng2024llamafactory. Training is performed over two epochs on approximately 17k examples, which are presented sequentially, without shuffling, reflecting a curriculum of increasing logical complexity. Training is distributed across 8 GPUs using DeepSpeed with ZeRO Stage 3 optimization, taking 4 hours. Both optimizer states and model parameters are offloaded to CPU with pinned memory to maximize GPU memory efficiency. The AdamW optimizer is used in conjunction with a Warmup Cosine learning rate scheduler. Mixed-precision training is employed, with both bfloat16 and fp16 enabled in automatic mode. Communication overlap and contiguous gradients are activated to improve throughput, and model weights are saved in 16-bit precision at each checkpoint. Due to memory limitations, input sequences are truncated to a maximum length of 6k tokens using the Llama3 template, restricting training to slr_basic_train, slr_easy_train, and slr_medium_train splits. Optimization is performed using cross-entropy loss over the ground truth rule $R^{\star}$, with a per-device batch size of 5 and gradient accumulation over 2 steps, resulting in an effective batch size of 80 samples per step across 8 GPUs. The learning rate is set to $2\times 10^{-4}$, scheduled with a cosine scheduler and a warmup ratio of 0.03. All relevant hyperparameters and training scripts are included in the codebase for full reproducibility.
G.2 Evaluation Setup
For downstream evaluation, we use the Language Model Evaluation Harness evalharness with default settings for each benchmark, enabling few-shot as multiturn prompting to support multi-turn contexts where applicable, including the official pass rate (typically pass@1) and whether evaluation is performed in zero-shot or few-shot mode. All evaluations are conducted on 8 GPUs using vLLM kwon2023efficient for efficient batch inference. Reported scores reflect accuracy for each model and benchmark, and all results are based on the official evaluation splits and standardized prompt formatting consistent with the SLR curriculum.
Appendix H Code and Licenses
Appendix I Potential Risks

### v5

@@ Appendix C Compute Costs and Model Pricing
Compute costs are reported as the total USD cost to run all prompts, based on publicly listed API prices as of 01.05.2025, with Qwen3 and GPT5 model pricing added on 01.12.2025. Pricing ignores server-side token caching, as actual cache hit counts are unavailable. Table 5 summarizes per-model cost rates and API sources.
⟦t⟧ Dataset Scope & Task Reasoning Specialty Reference MMLU Multi-domain analytical reasoning Factual knowledge, mathematics, statistics, computer science hendryckstest2021 CLUTRR Relational and multi-hop inference Structured family graphs sinha2019clutrr LogiQA (2.0) Formal deductive reasoning Argument consistency liu2020logiqa; Hanmeng2023logicqa2 GPQA High-level scientific reasoning Graduate-level difficulty rein2024gpqa ARC Symbolic abstraction Pattern-based problem solving Clark2018ThinkYH HellaSwag Grounded commonsense inference Contextual continuity zellers2019hellaswag
Table 6: Downstream reasoning datasets. Overview of downstream evaluation datasets and what they evaluate for.
Appendix D Training and Evaluation Details
D.1 Training Setup
For our experiments on SLR-Bench, we fine-tune Llama-3.1-8B-Instruct using supervised fine-tuning (SFT) with LoRA adapters using LLaMA-Factory zheng2024llamafactory. Training is performed stage-wise, presenting basic, easy, and medium problems sequentially, reflecting a curriculum of increasing logical complexity. Training is distributed across 8 GPUs using DeepSpeed with ZeRO Stage 3 optimization, taking 4 hours. Both optimizer states and model parameters are offloaded to CPU with pinned memory to maximize GPU memory efficiency. The AdamW optimizer is used in conjunction with a Warmup Cosine learning rate scheduler. Mixed-precision training is employed, with both bfloat16 and fp16 enabled in automatic mode. Communication overlap and contiguous gradients are activated to improve throughput, and model weights are saved in 16-bit precision at each checkpoint. Due to memory limitations, input sequences are truncated to a maximum length of 6k tokens using the Llama3 template, restricting training to slr_basic_train, slr_easy_train, and slr_medium_train splits. Optimization is performed using cross-entropy loss over the ground truth rule $R^{\star}$, with a per-device batch size of 5 and gradient accumulation over 2 steps, resulting in an effective batch size of 80 samples per step across 8 GPUs. The learning rate is set to $2\times 10^{-4}$, scheduled with a cosine scheduler and a warmup ratio of 0.03. All relevant hyperparameters and training scripts are included in the codebase for full reproducibility.
D.2 Evaluation Setup
To evaluate the generality of reasoning skills induced by SLR, we benchmark models on a diverse set of widely used reasoning datasets. These benchmarks span multiple reasoning modalities, including multi-domain analytical reasoning, relational and multi-hop inference, formal deductive reasoning, symbolic abstraction, scientific reasoning, and grounded commonsense inference. Together, they provide broad coverage of the reasoning capabilities relevant for modern large language models. An overview of all downstream datasets, including their scope, reasoning focus, and references, is provided in App. Tab. 6. MMLU and its subsets assess analytical reasoning across factual knowledge, mathematics, statistics, and computer science hendryckstest2021. CLUTRR evaluates relational and multi-hop inference over structured family graphs sinha2019clutrr. LogiQA and LogiQA2 focus on formal deductive reasoning and logical consistency in natural language arguments liu2020logiqa; Hanmeng2023logicqa2. GPQA and its variants (Extended and Diamond) measure graduate-level scientific reasoning rein2024gpqa. ARC-Easy and ARC-Challenge test symbolic abstraction and pattern-based reasoning Clark2018ThinkYH, while HellaSwag evaluates grounded commonsense inference and contextual continuity zellers2019hellaswag. For downstream evaluation, we use the Language Model Evaluation Harness evalharness with the default protocol of each benchmark, enabling few-shot as multiturn prompting to support multi-turn contexts where applicable, including the official pass rate (typically pass@1) and whether evaluation is performed in zero-shot or few-shot mode. CLUTRR, which is not supported by evalharness, follows a similar setup with accuracy averaged across all test sets. All evaluations are conducted on 8 GPUs using vLLM kwon2023efficient for efficient batch inference. Reported scores reflect accuracy for each model and benchmark, and all results are based on the official evaluation splits and standardized prompt formatting consistent with the SLR curriculum.
Appendix E Ablation Studies
E.1 Ablation across synthesis parameters
⟦t⟧ Ablation Factor Configuration Accuracy (%) Rule Sampling Uniform 11.0 (↓9.5) LLM-guided 2.5 (↓8.6) Problem Size ($\boldsymbol{\kappa}$) $\kappa=4$ 35.0 (↓9.5) $\kappa=6$ 30.0 (↓5.0) ($\boldsymbol{R_{\text{len}}}$) 1–2 literals 7.2 (↓9.5) 2–3 literals 2.6 (↓4.6) Background Mirror 30.0 (↓9.5) Sampling Uniform 20.5 (↓9.5)
Table 7: Ablations Across Synthesis Parameters. Accuracy (%) averaged across the LLaMA base models for different synthesis parameters. Each row corresponds to one ablation factor, comparing paired configurations. Arrows in red indicate decreases relative to the first configuration in each pair.
To assess how synthesis configurations affect task difficulty, we conducted controlled ablations across the rule sampling policy, problem size, rule length, and background sampling policy, averaging results across all open-source LLaMA base models. As summarized in App. Tab. 7, these experiments confirm that SLR yields interpretable and controllable gradients in logical complexity. For each factor, we selected representative levels of the dataset where the targeted parameter changes.
(1) Rule sampling policy. To measure the influence of rule generation, we compare subsets within the easy-difficulty tier (levels 6–10) containing either uniformly sampled versus LLM-guided rules. As shown in Table 7, accuracy drops sharply from 11.0 % $\rightarrow$ 2.5 %, demonstrating that LLM-guided synthesis introduces more structured and semantically complex rules that substantially increase reasoning difficulty.
(2) Problem size ($\boldsymbol{\kappa}$). To isolate the effect of larger reasoning tasks, we compare level 4 to level 5, where $\kappa$ increases from 4 to 6. This change leads to a moderate decrease in accuracy (35.0 % $\rightarrow$ 30.0 %), showing that models begin to struggle as the relational scope and combinatorial complexity expand.
(3) Rule length ($\boldsymbol{R_{\text{len}}}$). To examine the effect of rule depth, we compare level 8 (rules with one–two literals) and level 9 (two–three literals). Accuracy drops markedly from 7.2 % $\rightarrow$ 2.6 %, confirming that even modestly longer reasoning chains substantially raise logical complexity and error rates.
(4) Background sampling policy. Finally, we compare level 5, which employs mirror sampling, to level 6, which uses uniform sampling over the background facts. Accuracy decreases from 30.0 % $\rightarrow$ 20.5 %, indicating that the mirror trains are significantly easier to solve than the uniformly sampled trains.
Together, these ablations highlight that the synthesis parameters in SLR modulate task difficulty, confirming that SLR-Bench provides a precise and controllable testbed for analyzing logical reasoning performance across model families.
⟦t⟧ Curriculum LRL Basic Easy Medium Hard ($\uparrow$0-20) ($\uparrow$%) ($\uparrow$%) ($\uparrow$%) ($\uparrow$%) Ordered 6.01 88.7 28.4 3.3 0.2 Random 5.99 88.1 28.5 3.6 0.05 Reverse 5.93 86.8 27.7 4.5 0.05
Table 8: Curriculum Learning Order. Comparison of curriculum learning orders (ordered, random, and reverse) under matched training budgets (levels 5–7) using llama3.1-8b-it. The ordered curriculum achieves the highest overall LRL.
E.2 Curriculum Learning Order
To evaluate the effect of curriculum ordering on reasoning acquisition, we conducted a controlled ablation comparing standard curriculum learning, random order, and reverse curriculum training. All settings were matched for dataset, sample size, and total token budget, using the same difficulty levels (5–7). As shown in Tab. 8, the curriculum-trained model achieves the highest overall Logical Reasoning Level (LRL) and more balanced performance across tiers, while random and reverse orders yield slightly lower LRL and less stable behavior across difficulty levels. These results indicate that a structured progression of reasoning difficulty facilitates smoother learning and generalization in SLR.
⟦t⟧ Model Variant LRL Basic Easy Medium Hard Cost (\$) o4-mini-low 10.3 91 81 25 9 7.26 o4-mini 12.3 93 88 52 13 21.43 o4-mini-high 12.8 98 96 40 21 24.24
Table 9: Scaling Test-time Compute within the o4-mini Family. Increasing compute improves reasoning ability but raises inference costs and does not yield uniform accuracy gains across all tiers.
E.3 Scaling Test-time compute with o4-mini
Table 9 analyzes the impact of increasing test-time compute within the o4-mini family. The high variant achieves the highest overall LRL, outperforming both o4-mini and o4-mini-low, but incurs a cost increase of over 13%. Notably, it performs worse on the medium tier (40% vs. 52%), indicating diminishing returns from additional compute. This supports the main text finding that higher inference cost does not guarantee proportional reasoning gains.
E.4 Compute Increases with Task Complexity
⟦t!⟧ Refer to captionwidth=⟧figs/completion_tokens_vs_accuracy_combined.pdf
As reasoning tasks grow in complexity, models generally require more completion tokens to solve them, leading to higher inference costs (see Fig. 5). This trend serves as a sanity check confirming that SLR-Bench ’s task difficulty correlates with increased reasoning effort rather than being arbitrary. However, since larger models inherently consume more compute per token, these results should not be interpreted as direct efficiency comparisons across models.
⟦t⟧ LRL ($\uparrow$0-20) Syntax($\uparrow$%) Basic($\uparrow$%) Easy($\uparrow$%) Medium($\uparrow$%) Hard($\uparrow$%) Qwen3-8B 4.6 93 79 14 0 0 Qwen3-8B-SLR 7.9 (+3.3) 100 (+7) 96 (+17) 43 (+29) 14 (+14) 6 (+6) Downstream Reasoning Performance SLR-Homes ($\uparrow$%) MMLU ($\uparrow$%) MMLU-Stats ($\uparrow$%) CLUTRR ($\uparrow$%) LogiQA ($\uparrow$%) LogiQA2 ($\uparrow$%) Qwen3-8B 38.0 49.8 18.1 19.6 25.0 27.7 Qwen3-8B-SLR 40.7 (+2.7) 55.9 (+6.1) 18.1 (+0.0) 20.6 (+1.0) 31.5 (+6.5) 31.0 (+3.3) GPQA ($\uparrow$%) GPQA-Ext. ($\uparrow$%) GPQA-Dia. ($\uparrow$%) ARC-Easy ($\uparrow$%) ARC ($\uparrow$%) HellaSwag ($\uparrow$%) Qwen3-8B 23.4 26.4 21.2 69.2 43.3 48.8 Qwen3-8B-SLR 26.8 (+3.4) 27.1 (+0.7) 27.8 (+6.5) 77.3 (+8.1) 49.8 (+6.5) 51.6 (+2.8)
Table 10: Qwen3-8B-SLR. Benchmark scores ($\uparrow$) for base and SLR-tuned models on SLR-Bench and downstream benchmarks; LRL measuring cumulative curriculum progress. The tuned model surpasses the baseline across all curriculum stages, while generalizing to other downstream reasoning tasks.
Appendix F Additional Experiments
F.1 Extended Leaderboard
We extend our evaluation to the Qwen3 and GPT-5 model families and consolidate all new model results into a single unified extended leaderboard (Tab. 16). The leaderboard compares reasoning-oriented models (orange) with non-reasoning models (blue) across model families and scales, providing a comprehensive overview of current reasoning performance on SLR-Bench. The more recent models show improved LRL. Despite these advances, GPT-5 solves only 46% of the tasks in the Hard tier, highlighting that SLR-Bench remains challenging even for the strongest contemporary models, particularly at higher levels of inductive complexity.
F.2 Qwen3-SLR
We fine-tune a LoRA adapter on Qwen3-8B using a similar setup as the llama model, but with SLR training data shuffled. Subsequently, we evaluate the resulting model on SLR-Bench and on downstream reasoning benchmarks. As shown in App. Tab. 10, Qwen-SLR consistently improves performance on SLR tasks and also yields gains across downstream reasoning benchmarks, indicating that the learned reasoning capabilities transfer beyond the training distribution.
F.3 DeepSeek-SLR
⟦t⟧ LRL ($\uparrow$0-20) Syntax($\uparrow$%) Basic($\uparrow$%) Easy($\uparrow$%) Medium($\uparrow$%) Hard($\uparrow$%) DeepSeek-R1-1.5B 0 54 0 0 0 0 DeepSeek-1.5B-SLR 5.6 (+5.6) 97(+43) 84 (+84) 27 (+27) 2 (+2) 0 (+0)
Table 11: DeepSeek-SLR. GRPO training of DeepSeek-R1-1.5B on SLR-Bench. Benchmark scores ($\uparrow$%) for base and SLR-tuned models (DeepSeek-R1-1.5B) on SLR-Bench; LRL measures cumulative curriculum progress. The SLR-tuned model improves syntax and reasoning capabilities on basic, easy, and medium problems.
To examine whether the SLR curriculum remains effective in RL setups, we applied it using Group Reinforcement Preference Optimization (GRPO) Shao2024grpo on the small-scale reasoning model deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B. All training was performed under the same setup as SFT. The resulting model exhibits a clear improvement in reasoning ability, with the Logical Reasoning Level (LRL) increasing from 0.0 to 5.6. Performance gains are observed across reasoning tiers, especially on basic and easy reasoning tasks, demonstrating that a structured progression of task difficulty benefits even small models trained via reinforcement optimization. However, training was halted at level 7 due to reward collapse in later stages, suggesting that smaller architectures experience instability once the reasoning task complexity surpasses their capacity. Overall, these results indicate that the SLR curriculum not only supports SFT but also effectively enhances reasoning in RL setups (GRPO), reinforcing its flexibility as a general training paradigm.
F.4 Extended Pass@k and CoT Results
[…]
Table 12: SLR-Bench Pass@k and CoT Results. We report Logical Reasoning Level (LRL) scores at Pass@1, Pass@4, and Pass@8 (%) for open-source models, together with total completion tokens (M) and inference cost (\$). Higher LRL values indicate stronger logical reasoning on SLR-Bench. Chain-of-Thought (CoT) variants illustrate the effect of explicit reasoning traces, while increasing $k$ reflects test-time sampling improvements.
To complement the main results, where we report average Pass@1 (LRL) performance, we provide extended Pass@k scores (averaged for $k<8$) and Chain-of-Thought (CoT) results for open-source models in App. Tab. 12. The table summarizes model performance at different test-time sampling levels ($k=1,4,8$), together with compute requirements in tokens and cost for a single pass. Across all models, accuracy increases with larger $k$, confirming that multiple sampled completions raise the probability of generating a correct reasoning trace. Beyond improving performance, Pass@k also reveals reasoning variance. Smaller models (e.g., 3B–8B) show larger Pass@1–Pass@8 gaps, doubling in accuracy as we move from pass@1 to pass@8, indicating less stable reasoning, while larger models display narrower gaps and more consistent inference. CoT variants generally show modest gains over their non-CoT counterparts of comparable scale, indicating that explicit reasoning traces can improve reasoning, but also come with higher computational costs. Overall, these findings show that both test-time sampling and CoT prompting not only enhance reasoning outcomes but also serve as useful probes into a model’s internal reasoning diversity and reliability.
F.5 CLUTRR Evaluation
⟦t⟧ Llama-3.1-8B-it CLUTRR Variant Base w/ SLR w/ DeepSeek-R1 gen_train23_test2to10 10.2 19.1 24.0 gen_train234_test2to10 9.5 16.4 25.0 rob_train_clean_23_test_all 29.1 35.6 26.0 rob_train_sup_23_test_all 43.2 45.2 23.0 rob_train_irr_23_test_all 31.3 34.5 20.0 rob_train_disc_23_test_all 29.7 41.6 18.0 Average 25.5 32.1 23.0
Table 13: CLUTRR Evaluation. Accuracy (%) of base, SLR-tuned, and reasoning models on CLUTRR test sets (greedy decoding). SLR-tuned outperforms both base and reasoning models on CLUTRR.
We evaluated SLR-tuned and baseline models on the CLUTRR benchmark sinha2019clutrr, an inductive reasoning dataset requiring inference over kinship relations. All models were evaluated with greedy decoding, focusing on accuracy since CLUTRR uses a closed classification space where syntax validity does not apply. Across all variants, SLR-tuning consistently improves over the base model, with an average gain of +6.6pp, and remains competitive with or superior to DeepSeek-R1-8B on most configurations (App. Tab. 13). Notably, Llama-3.1-8B-SLR achieves higher accuracy than DeepSeek-R1-8B on four of six variants while using dramatically fewer resources—only 11k completion tokens vs. 8.1M for DeepSeek (730× less compute). These results provide strong evidence that SLR enhances inductive reasoning beyond SLR-Bench, achieving higher accuracy and far greater computational efficiency on an established external inductive reasoning benchmark.
Appendix G Error Mode Analysis.
To better understand model failures on SLR-Bench, we conducted a comprehensive error analysis of generated rules. As discussed in Section 4, syntax is rarely the limiting factor; most models achieve near-perfect syntactic validity. Consequently, the primary bottleneck lies in semantic rule induction.
G.1 Taxonomy of Error Modes
By inspecting a sample of failed generations across all models, we identified five prominent failure patterns:
  1. 1.
    Over-generalization: Rules that are too broad, often consisting of only one or two predicates (e.g., eastbound(T):has_car(T,C).eastbound(T):-has\_car(T,C).).
  2. 2.
    Under-generalization: Rules that are overly specific, often containing excessive literals (more than eight) that fail to capture the concise ground truth (e.g., $eastbound(T):-has\_car(T,C),car\_len(C,short),\dots$).
  3. 3.
    Failure of Abstraction: The model lacks semantic understanding of the induction task. Instead of lifting specific examples into variables, the model overfits to the prompt context, replacing general logic with grounded atoms (e.g., hardcoding eastbound(T):car1eastbound(T):-car1). This suggests the model is performing text completion on the prompt rather than logical reasoning.
  4. 4.
    Degenerate Reasoning: The model falls into repetition loops, meaningless enumerations (e.g., repeating “iiii”), or fails to produce structured logic entirely.
  5. 5.
    Shortcut behavior (Reward Hacking): The model completely disregards the task’s objective and attempts to take a shortcut. Instead of finding a single universally quantified rule that generalizes the underlying patterns, the model tries to take a shortcut and directly assign the labels (e.g., eastbound(train1), eastbound(train2), …).
G.2 Quantitative analysis
Quantitative analysis across all models reveals that 7% of generations were short rules ($\leq$2 literals, 38% success), 20% overly long rules ($\geq$8 literals, 14% success), 1% had grounded heads (0% success), and 8% showed degenerate patterns (33% success). Degenerations were identified via simple heuristics (e.g., $\geq 10$ repeated characters, $\geq$10 enumerated lines, or $\geq 30\%$ dominance of a single character). While approximate, these checks capture systematic breakdowns. Notably, even degenerate outputs sometimes partially recover, with 25% still solving the task.
G.3 Model-Specific Behaviors
While most failure modes emerge across all models, some failure modes are predominantly assignable to some model classes.
GPT-5 Shortcut Behavior.
The strongest reasoning models (GPT-5, GPT-5-mini) occasionally exhibit behavior consistent with reward hacking. Instead of inducing a general logic rule, these models attempt to "bypass" the task by directly assigning labels to the constants found in the prompt (Error Type 5).
⟦t⟧ # Shortcut Model Basic Easy Medium Hard $\sum$ gpt-5 0 0 3 1 4 gpt-5-mini-high 0 1 11 28 40 gpt-5-mini 0 0 8 8 16 gpt-5-mini-low 0 0 0 0 0 gpt-5-nano 0 2 6 0 8 gpt-4.5-preview 0 0 0 0 0 gpt-4o 0 0 0 0 0 gpt-4-turbo 0 0 0 0 0
Table 14: GPT-5 Shortcut Analysis. Number of Shortcuts attempted in each complexity tier. Non-reasoning models do not attempt shortcuts. In reasoning models, the frequency of attempted shortcuts increases with both task complexity and the model’s reasoning effort.
As shown in Table 14, this behavior is not random. The gpt-5-mini-high variant attempts a total of 40 such shortcuts, with the frequency increasing alongside task complexity. A plausible explanation is rooted in RL-based training objectives: if training rewards insufficiently penalize grounded hypotheses, or if proxy rewards correlate imperfectly with true semantic understanding, models may exploit these loopholes to maximize apparent reward. Our symbolic judge correctly rejects these as unsolved.
Degenerate Reasoning in Distilled Reasoning Models.
In contrast, distilled reasoning models (e.g., DeepSeek-R1-Llama-70B) struggle with Degenerate Reasoning (Error Type 4). Unlike smaller models that simply fail to abstract, these larger distilled models attempt complex reasoning but frequently derail.
⟦t⟧ Decile (short $\rightarrow$ long) 0 1 2 3 4 5 6 7 8 9 Accuracy ($\uparrow$) 96 87 87 80 42 7 2 0 0 0
Table 15: Success Rate by Output Length Decile. Binning generations by relative output length (short $\rightarrow$ long) on DeepSeek-R1-Llama-80B. Longer reasoning traces correlate strongly with lower accuracy, indicating that verbosity often signals derailment rather than deeper reasoning.
As detailed in Table 15, we observed a strong inverse correlation between reasoning length and success. Success rates drop from 96% in the shortest decile to 0% in the longest, indicating verbosity reflects derailment rather than deeper reasoning.
Failure of Abstraction in Small Models.
Conversely, smaller distilled models (e.g., Llama-3.2-1B) struggle with fundamental Failure of Abstraction (Error Type 3). These models lack the capacity to maintain the distinction between the example data (constants) and the logical rule (variables). They rarely attempt shortcuts or fall into complex reasoning loops; instead, they simply parrot the input context or hallucinate grounded atoms, effectively treating the logic induction task as a text completion exercise.
Appendix H Code and Licenses
Appendix I Potential Risks
@@ 3. Instruction Prompt (example formats):
    % Find a rule "is_red_train(T) :-" that solves the bk.
@@
    % Find a rule "is_red_train(T) :-" that solves the bk.
@@
[…]
Table 16: Extended Leaderboard. Logical Reasoning Level (LRL), syntax score, and stage-specific logical reasoning accuracy (basic, easy, medium, hard) and computational costs. Models are ranked by LRL. Models with reasoning mode enabled outperform their base counterparts across reasoning stages, though performance decreases as task complexity rises.